# PhishDebate: An LLM-Based Multi-Agent Framework for Phishing Website Detection

### v1

@@ Abstract
Phishing websites continue to pose a significant cybersecurity threat, often leveraging deceptive structures, brand impersonation, and social engineering tactics to evade detection. While recent advances in large language models (LLMs) have enabled improved phishing detection through contextual understanding, most existing approaches rely on single-agent classification facing the risks of hallucination and lack interpretability or robustness. To address these limitations, we propose PhishDebate, a modular multi-agent LLM-based debate framework for phishing website detection. PhishDebate employs four specialized agents to independently analyze different textual aspects of a webpage—URL structure, HTML composition, semantic content, and brand impersonation—under the coordination of a Moderator and a final Judge. Through structured debate and divergent thinking, the framework delivers more accurate and interpretable decisions. Extensive evaluations on commercial LLMs demonstrate that PhishDebate achieves 98.2% recall and 98.2% True Positive Rate (TPR) on a real-world phishing dataset, and outperforms single-agent and Chain of Thought (CoT) baselines. Additionally, its modular design allows agent-level configurability, enabling adaptation to varying resource and application requirements.
Index Terms:
Phishing Website Detection, Large Language Model, Multi-Agent System.
@@ I Introduction
Phishing attacks continue to pose a pervasive and evolving threat across the digital landscape, targeting individuals, businesses, and institutions alike [1]. By mimicking legitimate websites and exploiting human trust, phishing websites deceive users into disclosing sensitive information such as login credentials, financial details, or personal data [2, 3]. The sophistication and scale of phishing operations have escalated in recent years, often leveraging automation, obfuscation, and brand impersonation to evade detection systems [4]. Consequently, phishing remains one of the most pressing cybersecurity challenges, accounting for a significant proportion of fraud and data breaches [5].
To combat this threat, a wide range of detection techniques have been proposed. Traditional approaches relied on heuristic rules and blacklists, which are lightweight but prone to obsolescence in the face of adaptive attackers [6, 7]. Machine learning models have since emerged, incorporating engineered features from URLs, HTML structures, and third-party metadata [8, 9]. More recently, deep learning and transformer-based architectures have demonstrated promising results by learning patterns from raw inputs such as URLs, webpage content, and screenshots [10, 11, 12]. With the rise of Large Language Models (LLMs), phishing detection has entered a new era, where contextual understanding and semantic reasoning allow for the identification of subtle social engineering tactics [13].
However, existing LLM-based methods for phishing website detection rely on single-agent classification [14], which not only lacks iterative reasoning and interpretability but also risks hallucinations [15] and narrow decision-making due to dependence on a single perspective.
To address these limitations, we propose PhishDebate — a modular multi-agent LLM-based debate framework for phishing website detection. PhishDebate comprises four specialized agents, each targeting a distinct dimension of phishing evidence, including URL structure, HTML composition, semantic content, and brand impersonation. A Moderator oversees the structured debate process, while a final Judge delivers the classification verdict based on all arguments. The framework follows a modular design that allows users to flexibly include or exclude agents based on resource constraints or task-specific requirements. This structure supports collaborative reasoning, improves detection accuracy, enhances interpretability, and reduces dependence on a single-agent perspective.
@@
  • We introduce PhishDebate, one of the first debate-based multi-agent LLM frameworks for phishing website detection, emphasizing modularity, explainability, and collaborative inference.
@@
  • We conduct comprehensive evaluations using two real-world phishing datasets and four state-of-the-art LLMs, comparing PhishDebate against baselines including direct prompting and Chain of Thought (CoT) reasoning.
  • We demonstrate through scenario analysis and a case study that PhishDebate not only improves classification performance but also provides interpretable rationales, enhancing trust and transparency in phishing detection.
The remainder of this paper is organized as follows: Section II reviews related works in phishing detection and multi-agent LLM reasoning. Section III details the design of the PhishDebate framework, including agent roles and debate logic. Section IV presents the evaluation setup, datasets, and performance metrics. Section V reports experimental results, including comparative baselines and scenario analyses and a detailed case study. Finally, Section VI concludes the paper and outlines directions for future research.
II Related Works
@@ II-A Phishing Website Detection
Deep learning methods, especially Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs), further advanced detection capabilities by automatically learning patterns from raw input such as URLs, text, and screenshots [10, 11, 12]. Vision-based models using webpage screenshots and hybrid models combining URL, DOM, and content-level features have shown promise in detecting visually deceptive phishing websites [17, 18]. Recently, transformer-based architectures and pre-trained language models have been applied to capture semantic cues in webpage content and URLs [19, 1], improving the understanding of social engineering cues.
@@ III PhishDebate Framework
In this section, we present the detailed design of proposed PhishDebate framework and introduce specialist agents and coordination agents involved in this framework.
@@ III-A Overall PhishDebate Framework
Input: URL $U$, HTML $H$, text $T$, $R_{max}$, $\tau$, agents $A$
@@
             $context\leftarrow$ AggResp($responses$);
@@
       $decision\leftarrow$ QueryJudge($judge\_prompt$);
       return $(decision.assess,decision.conf,round,consensus.reached)$;
@@ III-B2 HTML Structure Agent
The HTML Structure Agent analyzes the underlying HTML code structure to identify technical indicators of phishing websites. This agent examines form elements, JavaScript usage, iframe implementations, and other structural patterns that may indicate malicious intent. The prompt template used for HTML Structure Agent can be found in supplementary materials [28].
@@ III-B3 Content Semantic Agent
The Content Semantic Agent performs natural language analysis of the visible website content to detect linguistic patterns and semantic cues associated with phishing attacks. This agent specializes in identifying manipulative language, urgency tactics, and social engineering techniques. The prompt template used for Content Semantic Agent can be found in supplementary materials [28].
@@ III-B4 Brand Impersonation Agent
The Brand Impersonation Agent specializes in detecting attempts to impersonate legitimate brands or organizations. This agent analyzes both URL and content elements to identify unauthorized use of brand names, logos, and corporate identity elements. The prompt template used for Brand Impersonation Agent can be found in supplementary materials [28].
@@ III-C1 Moderator
The Moderator serves as the debate coordinator, evaluating specialist agent responses after each round to determine if consensus has been reached. The Moderator analyzes the collective insights from all active specialist agents and decides whether the evidence is sufficient for a confident determination. The prompt template used for Moderator can be found in supplementary materials [28].
@@ III-C2 Judge
The Judge makes the final determination after the maximum number of debate rounds. The Judge considers all evidence presented by specialist agents throughout the entire debate process and renders a definitive verdict. The prompt template used for Judge can be found in supplementary materials [28].
@@
Moreover, PhishDebate follows a modular design philosophy that enables flexible configuration based on specific use cases and requirements. The framework supports several customization options:
@@ IV Evaluation
In this section, we introduce the LLMs and datasets used in this study and illustrate the data processing and evaluation metrics used.
IV-A Model Used
To evaluate the performance of the proposed PhishDebate framework, we selected four commercial LLMs for comparison and further analysis: Qwen2.5-vl-72b-instruct[29], Gemini-2.0-Flash[30], GPT-4o[31], and GPT-4o Mini[31]. These models were chosen for their advanced reasoning capabilities, diverse architectures, and wide availability via API, providing a representative benchmark across different commercial LLM ecosystems.
IV-B Dataset & data Processing
@@
We used the Mendeley phishing website dataset [32] for performance benchmarking and comparison across models, while both Mendeley [32] and TR-OP [37] datasets were used in the scenario analysis to enhance result robustness.
@@ IV-C Evaluation Metrics
[…]
@@ V-A Performance Evaluation
Table I illustrates the results of performance evaluation across four commercial LLMs on the PhishDebate framework, including GPT-4o, GPT-4o-mini, Gemini-2.0, and Qwen2.5-vl-72b-instruct. The results demonstrate that GPT-4o consistently outperforms others in terms of overall classification effectiveness, achieving the highest accuracy (96.50%), precision (94.97%), and F1 score (96.56%), along with the best TNR (94.8%) and the lowest FPR (5.2%). Gemini-2.0, while exhibiting the fastest inference time (17.9s), achieved the highest TPR (99.2%) and the lowest FNR (0.8%), indicating strong phishing detection capability. However, its performance was affected by 8 instances in which it failed to follow the binary classification instruction in the prompt, instead outputting an uncertain response. These cases, originating from legitimate websites, were treated as misclassifications, thereby increasing Gemini’s FPR to 14.4%, lowering its TNR to 85.57%, and reducing its overall accuracy and F1 score to 92.44% and 92.97%, respectively. GPT-4o-mini performs competitively but lags slightly behind GPT-4o, whereas Qwen showed the lowest accuracy and F1 score due to a significantly higher FNR (19.4%), despite maintaining a relatively low FPR (9.4%). These results highlight GPT-4o as the most balanced and reliable model for phishing detection in the context of the PhishDebate framework.
@@ V-B Comparative Evaluation
To evaluate the effectiveness of our proposed multi-agent debate system, we compare it against two widely adopted single-agent prompting strategies. First, we include a Single-Agent with Direct Prompt baseline, where a LLM receives the raw input (URL and HTML-visible text) and directly predicts whether the page is phishing without any intermediate reasoning. This baseline reflects a straightforward application of LLMs for classification tasks and serves as a reference point for measuring improvements attributable to structured reasoning. Second, we include a Single-Agent with CoT baseline [39], where the LLM is prompted to reason step-by-step through URL analysis, content interpretation, and brand impersonation detection before arriving at a decision. This variant captures the benefits of explicit intermediate reasoning within a single model, enabling us to isolate and assess the added value of distributed agent roles and debate-based aggregation in our system. We provide the prompt templates used for two baselines in supplementary materials [28].
Table. II presents the results of comparative evaluation between the proposed PhishDebate framework and two baselines—Single Agent and CoT prompting. The results demonstrate the effectiveness of our debate-driven approach in enhancing phishing detection on GPT-4o-mini. PhishDebate outperforms both baselines across all key evaluation metrics, achieving the highest precision (90.57%), accuracy (93.90%), recall (98.00%), and F1 score (94.14%). Notably, while CoT prompting provides reasonably strong results (e.g., 90.94% F1 score), it suffers from lower recall and precision compared to PhishDebate, especially when factoring in the 50 cases it labeled as “uncertain” instead of providing definitive classifications. These include 16 phishing websites and 34 legitimate websites, which, according to our evaluation protocol, are considered misclassifications. This underscores a critical limitation of CoT in reliability and confidence. In contrast, the PhishDebate framework significantly improves the confidence and decisiveness of LLM-based predictions by leveraging multi-agent debate, thereby reducing indecisive outputs and promoting consistent judgment across phishing and legitimate classes.
@@
Method Precision Accuracy Recall F1 Score
Single Agent 0.6057 0.6700 0.9740 0.7469
CoT 0.8861 0.9070 0.9340 0.9094
PhishDebate (Ours) 0.9057↑ 0.9390↑ 0.9800↑ 0.9414↑
@@ V-C Scenario Analysis
Table III presents a scenario analysis evaluating the performance of the PhishDebate framework when individual specialized agents are excluded. The results are averaged across two datasets (Mendeley [32] and TR-OP [37]) to ensure robustness. The full configuration using all agents achieves strong overall performance, particularly demonstrating the highest recall (0.985) and lowest FNR (FN = 8), indicating its superior ability to identify phishing websites with minimal missed detections. This highlights the importance of the ensemble debate mechanism in capturing diverse phishing indicators across URL, HTML, semantic, and brand features.
Notably, the exclusion of the HTML Structure Agent leads to the highest F1 score (0.9471), accuracy (0.9455), precision (0.9207), and the lowest FPR (FP = 42), suggesting that in specific deployment scenarios where legitimate website preservation is critical, a reduced configuration excluding this agent may offer better precision-legitimacy trade-offs. However, this comes at a minor cost in recall (0.975), which is still lower than the full configuration.
The variant without the Content Semantic Agent matches the full agent setup in recall and TPR (TP = 492), but suffers in terms of increased false positives (FP = 56), suggesting a reduced ability to discriminate legitimate content. This supports the value of semantic reasoning in reducing over-flagging. Similarly, excluding the URL Analyst Agent leads to a clear degradation in phishing detection capability, with a recall drop to 0.952 and increased false negatives (FN = 24), reinforcing the foundational role of URL features in phishing identification. Finally, omitting the Brand Impersonation Agent results in slightly lower overall performance, underscoring its contribution to brand spoof detection. Overall, these findings emphasize the effectiveness of the modular PhishDebate design, while demonstrating that task-specific configurations can still perform competitively under constrained scenarios.

### v2

@@ Abstract
Phishing websites remain a major cybersecurity threat, exploiting deceptive structures, brand impersonation, and social engineering to evade detection. Recent advances in large language models (LLMs) have improved phishing detection through contextual understanding, yet most existing approaches rely on single-agent classification, which is prone to hallucination and often lacks interpretability and robustness. To address these limitations, we propose PhishDebate, a modular multi-agent LLM-based debate framework for phishing website detection. Four specialized agents independently analyze webpage aspects, including URL structure, HTML composition, semantic content, and brand impersonation, under the coordination of a Moderator and final Judge. Through structured debate and divergent reasoning, the framework achieves more accurate and interpretable decisions. By reducing uncertain predictions and providing transparent reasoning, PhishDebate functions as an analyst-augmentation system that lowers cognitive load and supports early, left-of-exploit detection of phishing threats. Evaluations on commercial LLMs show that PhishDebate achieves 98.2% recall on a real-world phishing dataset and outperforms single-agent and Chain-of-Thought (CoT) baselines. Its modular design enables agent-level configurability, allowing adaptation to varying resource and application requirements, and offers scalability to high-velocity, large-scale security data environments.
@@ I Introduction
Phishing attacks continue to pose a pervasive and evolving threat across the digital landscape, targeting individuals, businesses, and institutions alike [18]. By mimicking legitimate websites and exploiting human trust, phishing websites deceive users into disclosing sensitive information such as login credentials, financial details, or personal data [9, 22]. The sophistication and scale of phishing operations have escalated in recent years, often leveraging automation, obfuscation, and brand impersonation to evade detection systems [19]. Consequently, phishing remains one of the most pressing cybersecurity challenges, accounting for a significant proportion of online fraud and data breaches [29].
To combat this threat, a wide range of detection techniques have been proposed. Traditional approaches relied on heuristic rules and blacklists, which are lightweight but prone to obsolescence in the face of adaptive attackers [11, 28]. Machine learning models have since emerged, incorporating engineered features from Uniform Resource Locators (URLs), HyperText Markup Language (HTML) structures, and third-party metadata [34, 35]. More recently, deep learning and transformer-based architectures have demonstrated promising results by learning patterns from raw inputs such as URLs, webpage content, and screenshots [25, 1, 6]. With the rise of Large Language Models (LLMs), phishing detection has entered a new era, where contextual understanding and semantic reasoning allow for the identification of subtle social engineering tactics [10].
However, most existing LLM-based methods for phishing website detection rely on single-agent classification [15], which not only lacks iterative reasoning and interpretability but also risks hallucinations [24] and narrow decision-making due to dependence on a single perspective. This limitation reduces their robustness and reliability when confronting complex, ambiguous, or deceptive phishing tactics.
To address these limitations, we propose PhishDebate, a modular multi-agent LLM-based debate framework for phishing website detection. PhishDebate comprises four specialized agents, each targeting a distinct dimension of phishing evidence, including URL structure, HTML composition, semantic content, and brand impersonation. A Moderator oversees the structured debate process, while a final Judge delivers the classification verdict based on all arguments. The framework follows a modular design that allows users to flexibly include or exclude agents based on resource constraints or task-specific requirements. This structure supports collaborative reasoning, improves detection accuracy, enhances interpretability, and reduces dependence on a single-agent perspective. Moreover, by reducing uncertain predictions and providing transparent reasoning, PhishDebate acts as an analyst-augmentation system that helps lower the cognitive burden of security analysts. Its ability to identify phishing attempts at an early, left-of-exploit stage underscores its relevance for proactive defense, while the modular design offers scalability to high-velocity, large-scale security data environments.
@@
  • We introduce PhishDebate, one of the first debate-based multi-agent LLM frameworks specifically designed for phishing website detection, emphasizing modularity, explainability, and collaborative inference.
@@
  • We conduct comprehensive evaluations using two real-world phishing datasets and four state-of-the-art LLMs (e.g., GPT-4o, Gemini-2.0), comparing PhishDebate against baselines including direct prompting and Chain of Thought (CoT) reasoning.
  • We demonstrate through scenario analysis and case studies that PhishDebate not only improves classification performance but also provides interpretable rationales, enhancing trust and transparency in phishing detection.
The remainder of this paper is organized as follows: Section II reviews relevant literature in phishing detection and multi-agent LLM reasoning. Section III details the design of the PhishDebate framework, including agent roles and debate logic. Section IV presents the evaluation setup, datasets, and performance metrics. Section V reports experimental results, including comparative baselines and scenario analyses and a detailed case study. Finally, Section VI discusses the limitations and Section VII concludes the paper with future directions.
II Related Work
@@ II-A Phishing Website Detection
Deep learning methods, especially Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs), further advanced detection capabilities by automatically learning patterns from raw input such as URLs, text, and screenshots [25, 1, 6]. Vision-based models using webpage screenshots and hybrid models combining URL, Document Object Model (DOM), and content-level features have shown promise in detecting visually deceptive phishing websites [23, 26]. Recently, transformer-based architectures and pre-trained language models have been applied to capture semantic cues in webpage content and URLs [27, 18], improving the understanding of social engineering cues.
@@ III PhishDebate Framework
In this section, we present PhishDebate, a novel multi-agent LLM framework for phishing website detection. Our approach leverages the collaborative intelligence of specialized agents through structured debate to achieve robust and interpretable phishing detection.
@@ III-A Overall PhishDebate Framework
Input: URL $U$, HTML $H$, text $T$, $R_{max}$, $\tau$, agents AA
@@
     $context\leftarrow$ AggResp(responsesresponses);
@@
  $decision\leftarrow$ QueryJudge(judge_promptjudge\_prompt);
  return (decision.assess,decision.conf,round,consensus.reached)(decision.assess,decision.conf,round,consensus.reached);
@@ III-B2 HTML Structure Agent
The HTML Structure Agent analyzes the underlying HTML code structure to identify technical indicators of phishing websites. This agent examines form elements, JavaScript usage, iframe implementations, and other structural patterns that may indicate malicious intent. The Fig. 3 details the prompt template used for HTML Structure Agent.
- Confidence: [A score between 0 and 1]
- Evidence: [Relevant structural elements or tag patterns you found]
Figure 3: Prompt template for HTML Structure Agent.
@@ III-B3 Content Semantic Agent
The Content Semantic Agent performs natural language analysis of the visible website content to detect linguistic patterns and semantic cues associated with phishing attacks. This agent specializes in identifying manipulative language, urgency tactics, and social engineering techniques. The Fig. 4 illustrates the prompt template used for Content Semantic Agent.
- Claim: [Whether the page language seems phishing-related]
- Confidence: [A score between 0 and 1]
Figure 4: Prompt template for Content Semantic Agent.
@@ III-B4 Brand Impersonation Agent
The Brand Impersonation Agent specializes in detecting attempts to impersonate legitimate brands or organizations. This agent analyzes both URL and content elements to identify unauthorized use of brand names, logos, and corporate identity elements. The Fig. 5 demonstrates the prompt template used for Brand Impersonation Agent.
Visible Text:
[TRUNCATED_VISIBLE_TEXT]
Provide response in following format:
- Claim: [Does the content attempt to impersonate a known brand?]
- Confidence: [A score between 0 and 1]
- Evidence: [Name(s) of impersonated brands and supporting context]
Figure 5: Prompt template for Brand Impersonation Agent.
@@ III-C1 Moderator
The Moderator serves as the debate coordinator, evaluating specialist agent responses after each round to determine if consensus has been reached. The Moderator analyzes the collective insights from all active specialist agents and decides whether the evidence is sufficient for a confident determination. The prompt template used for Moderator can be found in Appendix A Fig. 7.
@@ III-C2 Judge
The Judge makes the final determination when no consensus is reached after the maximum number of debate rounds. The Judge considers all evidence presented by specialist agents throughout the entire debate process and renders a definitive verdict. The prompt template used for Judge can be found in Appendix A Fig. 8.
@@
PhishDebate follows a modular design philosophy that enables flexible configuration based on specific use cases and requirements. The framework supports several customization options:
@@ IV Evaluation
IV-A Models Used
To evaluate the performance of the proposed PhishDebate framework, we selected four commercial LLMs for comparison and further analysis: Qwen2.5-vl-72b-instruct[4], Gemini-2.0-Flash[8], GPT-4o[30], and GPT-4o Mini[30]. These models were chosen for their advanced reasoning capabilities, diverse architectures, and wide availability via Application Programming Interface (API), providing a representative benchmark across different commercial LLM ecosystems.
IV-B Datasets & data Processing
@@
We used the Mendeley dataset for performance benchmarking and comparison across models, while both Mendeley and TR-OP datasets were used in the scenario analysis to enhance result robustness.
@@ IV-C Evaluation Metrics
  • F1 Score: $\text{F1}=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ Harmonic mean of precision and recall; balances false positives and false negatives.
@@ V-A Performance Evaluation
Table I illustrates the results of performance evaluation across four commercial LLMs on the PhishDebate framework, including GPT-4o, GPT-4o-mini, Gemini-2.0, and Qwen2.5-vl-72b-instruct. The results demonstrate that GPT-4o consistently outperforms others in terms of overall classification effectiveness, achieving the highest accuracy (96.50%), precision (94.97%), and F1 score (96.56%), along with the best true negative rate (94.8%) and the lowest false positive rate (5.2%). Gemini-2.0, while exhibiting the fastest inference time (17.9 s), achieved the highest true positive rate (99.2%) and the lowest false negative rate (0.8%), indicating strong phishing detection capability. However, its performance was affected by 8 instances in which it failed to follow the binary classification instruction in the prompt, instead outputting an uncertain response. These cases, originating from legitimate websites, were treated as misclassifications, thereby increasing Gemini’s false positive rate to 14.4%, lowering its true negative rate to 85.57%, and reducing its overall accuracy and F1 score to 92.44% and 92.97%, respectively. GPT-4o-mini performs competitively but lags slightly behind GPT-4o, whereas Qwen showed the lowest accuracy and F1 score due to a significantly higher false negative rate (19.4%), despite maintaining a relatively low false positive rate (9.4%). These results highlight GPT-4o as the most balanced and reliable model for phishing detection in the context of the PhishDebate framework.
@@ V-B Comparative Evaluation
To evaluate the effectiveness of our proposed multi-agent debate system, we compare it against two widely adopted single-agent prompting strategies. First, we include a Single-Agent with Direct Prompt baseline, where an LLM receives the raw input (URL and HTML-visible text) and directly predicts whether the page is phishing without any intermediate reasoning. This baseline reflects a straightforward application of LLMs for classification tasks and serves as a reference point for measuring improvements attributable to structured reasoning. Second, we include a Single-Agent with CoT baseline [37], where the LLM is prompted to reason step-by-step through URL analysis, content interpretation, and brand impersonation detection before arriving at a decision. This variant captures the benefits of explicit intermediate reasoning within a single model, enabling us to isolate and assess the added value of distributed agent roles and debate-based aggregation in our system. We provide the prompt templates used for two baselines in Appendix A Fig. 9 and Fig. 10.
Table II presents the results of comparative evaluation between the proposed PhishDebate framework and two baselines which are Single Agent and CoT prompting. The results demonstrate the effectiveness of our debate-driven approach in enhancing phishing detection on GPT-4o-mini. PhishDebate outperforms both baselines across all key evaluation metrics, achieving the highest precision (90.57%), accuracy (93.90%), recall (98.00%), and F1 score (94.14%). Notably, while CoT prompting provides reasonably strong results (e.g., 90.94% F1 score), it suffers from lower recall and precision compared to PhishDebate, especially when factoring in the 50 cases it labeled as “uncertain” instead of providing definitive classifications. These include 16 phishing websites and 34 legitimate websites, which, according to our evaluation protocol, are considered misclassifications. This underscores a critical limitation of CoT in reliability and confidence. In contrast, the PhishDebate framework significantly improves the confidence and decisiveness of LLM-based predictions by leveraging multi-agent debate, thereby reducing indecisive outputs and promoting consistent judgment across phishing and legitimate classes.
In terms of efficiency, the three approaches exhibit clear differences in average inference time (Table II). The Single Agent baseline is the fastest, requiring only 4.7 s per sample. CoT increases the processing cost to 10.5 s. By contrast, PhishDebate achieves the strongest detection effectiveness but requires 37.5 s. Despite the higher latency and additional tokens consumed by multi-round debates, the overall cost of running PhishDebate remains modest: processing 1,000 samples incurred only about \$3.357 in inference expenses.
@@
Method Prec. Acc. Rec. F1 Time
Single Agent 0.6057 0.6700 0.9740 0.7469 4.7 s↓
CoT 0.8861 0.9070 0.9340 0.9094 10.5 s
PhishDebate 0.9057↑ 0.9390↑ 0.9800↑ 0.9414↑ 37.5 s
Note. Prec. = Precision; Acc. = Accuracy; Rec. = Recall; F1 = F1 Score; Time = Avg. Time (seconds).
@@ V-C Scenario Analysis
Table III presents a scenario analysis evaluating the performance of the PhishDebate framework when individual specialized agents are excluded. The results are averaged across two datasets (Mendeley and TR-OP) to ensure robustness. The full configuration using all agents achieves strong overall performance, particularly demonstrating the highest recall (0.985) and lowest false negative rate (FN = 8), indicating its superior ability to identify phishing websites with minimal missed detections. This highlights the importance of the ensemble debate mechanism in capturing diverse phishing indicators across URL, HTML, semantic, and brand features.
Notably, the exclusion of the HTML Structure Agent leads to the highest F1 score (0.9471), accuracy (0.9455), precision (0.9207), and the lowest false positive rate (FP = 42), suggesting that in specific deployment scenarios where legitimate website preservation is critical, a reduced configuration excluding this agent may offer better precision-legitimacy trade-offs. However, this comes at a minor cost in recall (0.975), which is still lower than the full configuration.
The variant without the Content Semantic Agent matches the full agent setup in recall and true positive rate (TP = 492), but suffers in terms of increased false positives (FP = 56), suggesting a reduced ability to discriminate legitimate content. This supports the value of semantic reasoning in reducing over-flagging. Similarly, excluding the URL Analyst Agent leads to a clear degradation in phishing detection capability, with a recall drop to 0.952 and increased false negatives (FN = 24), reinforcing the foundational role of URL features in phishing identification. Finally, omitting the Brand Impersonation Agent results in slightly lower overall performance, underscoring its contribution to brand spoof detection. Overall, these findings emphasize the effectiveness of the modular PhishDebate design, while demonstrating that task-specific configurations can still perform competitively under constrained scenarios.